eeeeeeeeeeeeeeee:

Sophia needs $47 to go on a field trip. She has saved $11.50. She earns $6.50 per hour cleaning her neighbor's garden, and she earns $5.25 per hour selling cakes at a store. Can Sophia go on the field trip if she works at the garden for 4 hours and at the store for 2 hours? (Use the inequality 6.50x + 5.25y + 11.50 ≥ 47.) Yes, because the total will be $48 Yes, because the total will be $52 No, because the total will be $42 No, because the total will be $39

CripQUEZZ:

Ok so um first set it up 6.50x + 5.25y + 11.50 ≥ 47 So X = 4 And Y = 2 So do the multiplication first 6.50 x 4 = 26 And then 5.25 x 2 = 10.5 So now it’s 26 + 10.5 + 11.50 ≥ 47 So now jus simplify to see if it’s enough 26 + 10.5 = 36.5 Then 36.5 + 11.50 = 47 So it’s a

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ Ok so um first set it up 6.50x + 5.25y + 11.50 ≥ 47 So X = 4 And Y = 2 So do the multiplication first 6.50 x 4 = 26 And then 5.25 x 2 = 10.5 So now it’s 26 + 10.5 + 11.50 ≥ 47 So now jus simplify to see if it’s enough 26 + 10.5 = 36.5 Then 36.5 + 11.50 = 47 So it’s a \(\color{#0cbb34}{\text{End of Quote}}\) I mean 48 not 47